When 0.231G of phosphorus reacts with chlorine to form phosphorus trichloride, in a coffee-cup calorimeter of heat capacity 216J/C, the temperature of the calorimeter rises by 11.06C.

a)Provide a thermochemical equation

b) Give the Enthalpy change, Change of H, for the reaction
Answer #1

a) Thermal equation: 2P(s) + 3 Cl2(g) ----> 2PCl3(l)    ;   DHrxn = -641.34 KJ

b) No of mol of phosphorus reacted = 0.231/31 = 0.00745 mol

heat released(q) = C*DT

C = heat capacity = 216J/C

DT = 11.06 C

q = 216*11.06 = 2389 joule

q = 2.389 kj

DHrxn = -2.389/0.00745 = -320.67 kj/mol

DHrxn =-320.67*2 = -641.34 kj